OpenStudy (anonymous):

How do I write this as a composite of simpler functions I(x)=e^sin x help please!!!!!!!

OpenStudy (anonymous):

In which context did u see this problem?

OpenStudy (anonymous):

is it that reqd.. let f(x) =e^x and g(x)=sinx then I(x)=f(g(x))